OpenStudy (anonymous):

Simplify the expression and eliminate any negative exponents, assuming that all letters denote positive numbers. (x^6 * y)^(5/6) over (divided by) (y^4)^(5/6)

OpenStudy (anonymous):

is it this: \(\huge \frac{(x^6y)^{5/6}}{(y^4)^{5/6}} \) ???

OpenStudy (anonymous):

yes

OpenStudy (helder_edwin):

use these properties: \[ \large (ab)^n=a^nb^n \] \[ \large \frac{a^m}{a^n}=a^{m-n}=\frac{1}{a^{n-m}} \]

OpenStudy (helder_edwin):

also \[ \large (a^n)^m=a^{n\times m} \]

OpenStudy (anonymous):

all right. thanks!

OpenStudy (helder_edwin):

first work on the numerator and tell me what u get

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Is this it? I know I don't need the one under x^15, but the equation maker thingy isn't working amazingly well. \[\frac{ x^\frac{ 15 }{ 1 } }{ y^\frac{ 15 }{ 2 } }\]

OpenStudy (anonymous):

I also have another question like this. I'll put it in a new question so I can give you another medal:P

OpenStudy (helder_edwin):

wait.

OpenStudy (anonymous):

ok

OpenStudy (helder_edwin):

u got \[ \large \frac{(x^6y)^{5/6}}{(y^4)^{5/6}}=\frac{(x^6)^{5/6}y^{5/6}}{(y^4)^{5/6}} \]

OpenStudy (helder_edwin):

ok?

OpenStudy (anonymous):

Yeah, that was my first step.

OpenStudy (helder_edwin):

\[ \large =\frac{x^{6\times5/6}}{1} \left(\frac{y}{y^4}\right)^{5/6} \]

OpenStudy (anonymous):

i already got the answer, btw... i posted it a little earlier. sorry if you didn't see it!

OpenStudy (helder_edwin):

bear with me

OpenStudy (anonymous):

ok.

OpenStudy (helder_edwin):

\[ \large =\frac{x^5}{1}\frac{1}{(y^3)^{5/6}}= \frac{x^5}{1}\frac{1}{y^{3\times5/6}}=\frac{x^5}{y^{5/2}} \]

OpenStudy (helder_edwin):

this is the answer

OpenStudy (anonymous):

No, that's not right.

OpenStudy (helder_edwin):

yes it is

OpenStudy (anonymous):

No, it's not! one moment.

OpenStudy (anonymous):

Starts from the middle in the picture, on the left. Its my work.

OpenStudy (helder_edwin):

you posted 5/6 not 5/2

OpenStudy (anonymous):

Oh god, I'm sorry! oops.